# IP-LLM: Semantic Space Informed Prompt Learning with LLM for Time Series Forecasting

### v1

@@ Appendix D Full Few-shot Forecasting Results
| Methods | IP-LLM | Time-LLM | OFA | iTransformer | Dlinear | PatchTST | TimesNet | FEDformer | Autoformer | Stationary | ETSformer | LightTS | |||||||||||||
| MSE | MAE | MSE | MAE | MSE | MAE | MSE | MAE | MSE | MAE | MSE | MAE | MSE | MAE | MSE | MAE | MSE | MAE | MSE | MAE | MSE | MAE | MSE | MAE | ||
| Weather | 96 | 0.167 | 0.218 | 0.172 | 0.263 | 0.175 | 0.230 | 0.264 | 0.307 | 0.184 | 0.242 | 0.171 | 0.224 | 0.207 | 0.253 | 0.229 | 0.309 | 0.227 | 0.299 | 0.215 | 0.252 | 0.218 | 0.295 | 0.230 | 0.285 |
| 192 | 0.225 | 0.273 | 0.224 | 0.271 | 0.227 | 0.276 | 0.284 | 0.326 | 0.228 | 0.283 | 0.230 | 0.277 | 0.272 | 0.307 | 0.265 | 0.317 | 0.278 | 0.333 | 0.290 | 0.307 | 0.294 | 0.331 | 0.274 | 0.323 | |
| 336 | 0.280 | 0.320 | 0.282 | 0.321 | 0.286 | 0.322 | 0.323 | 0.349 | 0.279 | 0.322 | 0.294 | 0.326 | 0.313 | 0.328 | 0.353 | 0.392 | 0.351 | 0.393 | 0.353 | 0.348 | 0.359 | 0.398 | 0.318 | 0.355 | |
| 720 | 0.359 | 0.371 | 0.366 | 0.381 | 0.366 | 0.379 | 0.366 | 0.375 | 0.364 | 0.388 | 0.384 | 0.387 | 0.400 | 0.385 | 0.391 | 0.394 | 0.387 | 0.389 | 0.452 | 0.407 | 0.461 | 0.461 | 0.401 | 0.418 | |
| Avg | 0.257 | 0.295 | 0.260 | 0.309 | 0.263 | 0.301 | 0.309 | 0.339 | 0.263 | 0.308 | 0.269 | 0.303 | 0.298 | 0.318 | 0.309 | 0.353 | 0.310 | 0.353 | 0.327 | 0.328 | 0.333 | 0.371 | 0.305 | 0.345 | |
| Electricity | 96 | 0.153 | 0.251 | 0.147 | 0.242 | 0.143 | 0.241 | 0.162 | 0.264 | 0.150 | 0.251 | 0.145 | 0.244 | 0.315 | 0.389 | 0.235 | 0.322 | 0.297 | 0.367 | 0.484 | 0.518 | 0.697 | 0.638 | 0.639 | 0.609 |
| 192 | 0.169 | 0.268 | 0.158 | 0.241 | 0.159 | 0.255 | 0.180 | 0.278 | 0.163 | 0.263 | 0.163 | 0.260 | 0.318 | 0.396 | 0.247 | 0.341 | 0.308 | 0.375 | 0.501 | 0.531 | 0.718 | 0.648 | 0.772 | 0.678 | |
| 336 | 0.183 | 0.281 | 0.178 | 0.277 | 0.179 | 0.274 | 0.207 | 0.305 | 0.175 | 0.278 | 0.183 | 0.281 | 0.340 | 0.415 | 0.267 | 0.356 | 0.354 | 0.411 | 0.574 | 0.578 | 0.758 | 0.667 | 0.901 | 0.745 | |
| 720 | 0.239 | 0.324 | 0.224 | 0.312 | 0.233 | 0.323 | 0.258 | 0.339 | 0.219 | 0.311 | 0.233 | 0.323 | 0.635 | 0.613 | 0.318 | 0.394 | 0.426 | 0.466 | 0.952 | 0.786 | 1.028 | 0.788 | 1.200 | 0.871 | |
| Avg | 0.186 | 0.281 | 0.179 | 0.268 | 0.178 | 0.273 | 0.201 | 0.296 | 0.176 | 0.275 | 0.181 | 0.277 | 0.402 | 0.453 | 0.266 | 0.353 | 0.346 | 0.404 | 0.627 | 0.603 | 0.800 | 0.685 | 0.878 | 0.725 | |
| Traffic | 96 | 0.410 | 0.289 | 0.414 | 0.291 | 0.419 | 0.298 | 0.431 | 0.312 | 0.427 | 0.304 | 0.404 | 0.286 | 0.854 | 0.492 | 0.670 | 0.421 | 0.795 | 0.481 | 1.468 | 0.821 | 1.643 | 0.855 | 1.157 | 0.636 |
| 192 | 0.415 | 0.295 | 0.419 | 0.291 | 0.434 | 0.305 | 0.456 | 0.326 | 0.447 | 0.315 | 0.412 | 0.294 | 0.894 | 0.517 | 0.653 | 0.405 | 0.837 | 0.503 | 1.509 | 0.838 | 1.856 | 0.928 | 1.688 | 0.848 | |
| 336 | 0.433 | 0.310 | 0.437 | 0.314 | 0.449 | 0.313 | 0.465 | 0.334 | 0.478 | 0.333 | 0.439 | 0.310 | 0.853 | 0.471 | 0.707 | 0.445 | 0.867 | 0.523 | 1.602 | 0.860 | 2.080 | 0.999 | 1.826 | 0.903 | |
@@ Appendix F Visualization
In this section, we provide the visualizations of the forecasting cases of IP-LLM on ETTm1, Electricity, and Weather datasets under the input-512-predict-96 setting. As shown in Figure 6, IP-LLM achieves exceptionally good forecasting results across various datasets.

### v2

@@ Appendix D Full Few-shot Forecasting Results
| Methods | IP-LLM | Time-LLM(G) | OFA | iTrans. | Dlinear | PatchTST | TimesNet | FEDformer | Autoformer | Stationary | ETSformer | LightTS | |||||||||||||
| Data. | Hori. | MSE | MAE | MSE | MAE | MSE | MAE | MSE | MAE | MSE | MAE | MSE | MAE | MSE | MAE | MSE | MAE | MSE | MAE | MSE | MAE | MSE | MAE | MSE | MAE |
| Weather | 96 | 0.175 | 0.228 | 0.176 | 0.230 | 0.175 | 0.230 | 0.264 | 0.307 | 0.184 | 0.242 | 0.171 | 0.224 | 0.207 | 0.253 | 0.229 | 0.309 | 0.227 | 0.299 | 0.215 | 0.252 | 0.218 | 0.295 | 0.230 | 0.285 |
| 192 | 0.225 | 0.271 | 0.226 | 0.275 | 0.227 | 0.276 | 0.284 | 0.326 | 0.228 | 0.283 | 0.230 | 0.277 | 0.272 | 0.307 | 0.265 | 0.317 | 0.278 | 0.333 | 0.290 | 0.307 | 0.294 | 0.331 | 0.274 | 0.323 | |
| 336 | 0.282 | 0.321 | 0.292 | 0.325 | 0.286 | 0.322 | 0.323 | 0.349 | 0.279 | 0.322 | 0.294 | 0.326 | 0.313 | 0.328 | 0.353 | 0.392 | 0.351 | 0.393 | 0.353 | 0.348 | 0.359 | 0.398 | 0.318 | 0.355 | |
| 720 | 0.361 | 0.371 | 0.364 | 0.375 | 0.366 | 0.379 | 0.366 | 0.375 | 0.364 | 0.388 | 0.384 | 0.387 | 0.400 | 0.385 | 0.391 | 0.394 | 0.387 | 0.389 | 0.452 | 0.407 | 0.461 | 0.461 | 0.401 | 0.418 | |
| Avg | 0.260 | 0.297 | 0.264 | 0.301 | 0.263 | 0.301 | 0.309 | 0.339 | 0.263 | 0.308 | 0.269 | 0.303 | 0.298 | 0.318 | 0.309 | 0.353 | 0.310 | 0.353 | 0.327 | 0.328 | 0.333 | 0.371 | 0.305 | 0.345 | |
| Electricity | 96 | 0.148 | 0.248 | 0.148 | 0.248 | 0.143 | 0.241 | 0.162 | 0.264 | 0.150 | 0.251 | 0.145 | 0.244 | 0.315 | 0.389 | 0.235 | 0.322 | 0.297 | 0.367 | 0.484 | 0.518 | 0.697 | 0.638 | 0.639 | 0.609 |
| 192 | 0.159 | 0.255 | 0.160 | 0.257 | 0.159 | 0.255 | 0.180 | 0.278 | 0.163 | 0.263 | 0.163 | 0.260 | 0.318 | 0.396 | 0.247 | 0.341 | 0.308 | 0.375 | 0.501 | 0.531 | 0.718 | 0.648 | 0.772 | 0.678 | |
| 336 | 0.175 | 0.271 | 0.183 | 0.282 | 0.179 | 0.274 | 0.207 | 0.305 | 0.175 | 0.278 | 0.183 | 0.281 | 0.340 | 0.415 | 0.267 | 0.356 | 0.354 | 0.411 | 0.574 | 0.578 | 0.758 | 0.667 | 0.901 | 0.745 | |
| 720 | 0.235 | 0.326 | 0.236 | 0.329 | 0.233 | 0.323 | 0.258 | 0.339 | 0.219 | 0.311 | 0.233 | 0.323 | 0.635 | 0.613 | 0.318 | 0.394 | 0.426 | 0.466 | 0.952 | 0.786 | 1.028 | 0.788 | 1.200 | 0.871 | |
| Avg | 0.179 | 0.275 | 0.181 | 0.279 | 0.178 | 0.273 | 0.201 | 0.296 | 0.176 | 0.275 | 0.181 | 0.277 | 0.402 | 0.453 | 0.266 | 0.353 | 0.346 | 0.404 | 0.627 | 0.603 | 0.800 | 0.685 | 0.878 | 0.725 | |
| Traffic | 96 | 0.410 | 0.288 | 0.414 | 0.293 | 0.419 | 0.298 | 0.431 | 0.312 | 0.427 | 0.304 | 0.404 | 0.286 | 0.854 | 0.492 | 0.670 | 0.421 | 0.795 | 0.481 | 1.468 | 0.821 | 1.643 | 0.855 | 1.157 | 0.636 |
| 192 | 0.416 | 0.298 | 0.419 | 0.300 | 0.434 | 0.305 | 0.456 | 0.326 | 0.447 | 0.315 | 0.412 | 0.294 | 0.894 | 0.517 | 0.653 | 0.405 | 0.837 | 0.503 | 1.509 | 0.838 | 1.856 | 0.928 | 1.688 | 0.848 | |
| 336 | 0.435 | 0.313 | 0.438 | 0.315 | 0.449 | 0.313 | 0.465 | 0.334 | 0.478 | 0.333 | 0.439 | 0.310 | 0.853 | 0.471 | 0.707 | 0.445 | 0.867 | 0.523 | 1.602 | 0.860 | 2.080 | 0.999 | 1.826 | 0.903 | |
@@ Appendix F Visualization
In this section, we provide the visualizations of the forecasting cases of IP-LLM on ETTm2, Electricity, and Weather datasets under the input-512-predict-96 setting. As shown in Figure 7, IP-LLM achieves exceptionally good forecasting results across various datasets.
| Ablation Setting | Long-term Forecasting | |||
| ETTh2-96 | ETTh2-192 | ETTm2-96 | ETTm2-192 | |
| w/o Prompt & Alignment and w/o Decomposition | 0.289 | 0.358 | 0.170 | 0.231 |
| w/ Prompt & Alignment and w/o Decomposition | 0.287 | 0.353 | 0.166 | 0.228 |
| w/ Prompt & Alignment and w/ Decomposition | 0.278 | 0.346 | 0.165 | 0.222 |